In a recent study, 50 males used a new weight-loss supplement, and 40 of them experienced weight loss after two weeks. In the same study, 100 females used the same supplement, and 86 of them experienced weight loss after two weeks.

Fill in the blanks of the statement below to make the statement the most reasonable possible.

The new weight-loss supplement was less effective on ?malesfemales in the study because

%

of them failed to lose weight after two weeks, whereas only

%

of the ?malesfemales failed to lose weight after two weeks.
Answer #1

Males who failed to experience weight loss = 10/50 = 20%

Females who failed to experience weight loss = 14/100 = 14%

Hence,

The correct fillers in given blanks will be:

Males; 20; 14; Females